The surface of the Sun is approximately 5, 500 K,
Answer #1

Using a formula, we have

\DeltaS = S2 - S1 = (Q / T2) - (Q / T1)

\DeltaS = [(6000 J) / (291 K) - (6000 J) / (5500 K)]

\DeltaS = (20836 / 1067) J/K

\DeltaS = 19.5 J/K